OpenStudy (anonymous):

Susan Marciano invested part of her 40,000 bonus in a fund that paid an 11% profit and invested the rest in stock that surffered a 3% loss. Find the amount of each investment if her overall net profit was $1180.

OpenStudy (anonymous):

thought I answered this already.... 11% * X + (-3%) * (40000-x) = $1180 goes to.. .11x - ,03(40000-x) = 1180 .11x - 1200 +.03x = 1180 .14x = 1180 + 1200 .14x = 2380 x= 2380/.14 x = 17,000 so first investment was $17,000 and second was $23,000

OpenStudy (anonymous):

Thank you so much!!